# Beyond the Request: Harnessing HTTP Response Headers for Cross-Browser Web Tracker Classification in an Imbalanced Setting

### v1

@@ 3. Approach
In this study, we address the research gap identified in our review of related research by evaluating the (cross-browser) performance of commonly used machine learning classifiers that are trained on HTTP/S response headers. In this section, we discuss our underlying research questions and present the approach we used to address them.
@@ 3.1. Preliminaries
A core principle of automated web tracker detection is the definition of what constitutes a tracker. The focus hereby lies on the tracking entities and not on the activity of tracking. Our definition of what constitutes a tracker or non-tracker is coupled to our ground truth to label each dataset. We define them as follows:
@@
In other words: if the remote hostname (i.e. the sender) of a response matches at least one rule in our aggregate list compiled from the EasyList and EasyPrivacy filter lists, we label the response as a tracker.
For the sake of simplicity, we will use the terms header and response header interchangeably to refer to HTTP/S response headers throughout the remainder of the paper, unless explicitly stated otherwise.
@@ 3.2. Research Questions
The main goal of this study is to analyze whether response headers are useful for automated web tracker detection. Based on the ML approaches of prior research and the observed areas to further extend our empirical understanding and detection of web trackers, we formulate the following research questions that guide us throughout this study:
  1. (1)
    What are the differences and similarities between web trackers and non-trackers with regard to their response headers?
  2. (2)
    Can trained classifiers detect web trackers with only response header features while achieving an F1-Score higher than 0.9 in an imbalanced setting?
  3. (3)
    How large is a classifier’s performance degradation when it is trained on one browser’s HTTP traffic data and then applied to other browsers?
To answer these research questions, we develop a semi-automated ML pipeline that transforms HTTP traffic datasets collected with T.EX (Raschke et al., 2019) and uses it to train several supervised binary classifiers. We use reliable data that was compiled by prior research to facilitate future comparisons between new classifiers.
To test the robustness and reliability of our classifiers, we train them on HTTP data from Chrome and apply them to different datasets with a variety of configurations, including other browsers (latitudinal comparison) as well as different time frames within the same browser (longitudinal comparison). Furthermore, past research observed an uneven distribution of trackers and non-trackers, which is why we train our classifiers with an imbalanced distribution, to assess how effective they perform with the likely true distribution.
@@
Table 1. Overview of the collected data sets from (zen, 2023), split by data sets as well as tracker (T) and non-trackers (NT).
Figure 1. This figure illustrates the usage headers across four different browsers. The left plot (a) displays the frequency distribution of headers for each browser, showing how often each header is present in the dataset. The right plot (b) presents a boxen plot of the same data, which provides a more detailed view of the distribution, revealing variations and trends among the browsers.
@@ 3.3. Data Collection
We used existing data sets, containing labeled HTTP response headers by EasyList and EasyPrivacy. Following suggested good practices by (Demir et al., 2022), the crawl should fulfill three citeria: (i) repeated crawls across multiple browsers, namely Chrome, Firefox, and Brave, (ii) the documentation of the geographical region as well as the technical environment, and (iii) web sites from the Tranco ranking as it proved to be less biased than the Alexa ranking (Pochat et al., 2018). A previously crawled data set by Raschke using T.EX - The Transparency EXtension, fulfills all three criteria and is described as follows:
@@
After completing the crawls, each instance’s extension storage was extracted […]. Note: 2 of the 6 Firefox instances crashed during the crawl. Therefore, only 4 Firefox datasets are available (zen, 2023).
To assess a potential concept drift and therefore model degradation over time, we crawled one more data set from 2023 for Chrome. We followed a similar methodology as described in (zen, 2023), using T.EX with Chrome to explore the top 10K websites from the Tranco ranking as of March 29th, 2023.
Our reasons to choose only Chrome for a new data collection are as follows: Chrome is the most widely used in the real world with a market share of 66.13% on desktops, whereas Firefox has 7.1% (sta, 2023) and Brave has less than 1% (kin, 2023). Based on this distribution, the Chrome data set serves as the foundation for training to reflect and reach the majority, i.e. using data that most users encounter themselves. In addition, web privacy research often relies on a simulated Firefox browser with Selenium due to the use of OpenWPM, leaving space for other browsers of interest. In addition, it helps us to answer the question whether a model based on Chrome data can be applied to Firefox or Brave while achieving high performance metrics?.
@@ 4. Initial Response Header Analysis
To get a better understanding of the data underlying our analysis before designing our machine learning pipeline, we conduct an initial analysis of the data, the results of which we present in this section. We first explore response headers across all three browsers to find empirical evidence of their potential relation to trackers. Furthermore, we identify the characteristics of these datasets to apply the appropriate methods in our ML pipeline. We use these insights to develop our classifiers to evaluate the effectiveness of response headers for web tracker detection.
@@ 4.1. Descriptive Analysis of Headers
A descriptive analysis is useful to gain an understanding of the data structure and the distribution of classes and features within the dataset. Table 1 summarizes descriptive metrics for each dataset. We start with the datasets $\text{Chrome}_{22}$, $\text{Chrome}_{23}$, and $\text{Firefox}_{22}$. All three datasets show similar results for the distribution between trackers and non-trackers, the number of unique headers in total and per response. For example, we can observe a small quantitative change in the number of unique headers between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$. $\text{Chrome}_{22}$ lists a total of 7577 unique headers, while $\text{Chrome}_{23}$ has a slightly lower total of 7398 for non-trackers. When comparing trackers, $\text{Chrome}_{22}$ had 2213 unique headers, which is nearly the same as $\text{Chrome}_{23}$’s count of 2164. On average, non-trackers have more unique headers as well as a higher median $\tilde{\mu}$ and interquartile range than trackers.
@@
The $\text{Brave}_{22}$ browser with its standard configuration is the only one that has a low number of trackers and unique headers in tracking-related response messages. Interestingly, the $\tilde{\mu}$ equals 18 for trackers and 15 for non-trackers, which represents the only instance where trackers had a higher number. Although the dataset has almost only non-trackers, the number of unique headers is similar to the other datasets.
@@
Figure 1 shows a frequency plot, which is a distribution graph that shows how often each header appears in the dataset. It uses a logarithmic scale to manage the wide range of frequencies, which is a common approach when data spans several orders of magnitude. This plot typically demonstrates a power-law distribution, characterized by a steep drop-off where a small number of headers are occurring frequently, while the majority are rare. This type of distribution is common in HTTP data, where few headers dominate, and many others have a long tail of low frequency. This sparseness indicates that while there are many different headers, only a few are consistently used across different websites and by different browsers.
The boxen plot graph of Figure 1 offers a visual representation of the distribution of header frequencies across different browsers and versions. The spread and median values suggest that the distribution of headers is not uniform across browsers, with some browsers having a higher tendency to use certain headers than others. Furthermore, if the median line in the boxen plot is closer to the bottom of the box, it suggests a right-skewed distribution, aligning with what the frequency plot shows — a few headers with high frequency and many with low frequency.
@@
Figure 2. The Venn diagram reveals that a core set of 5672 headers are common across Chrome, Firefox, and Brave browsers across two years. Unique headers are present in each browser, with $\text{Chrome}_{23}$ having the highest number of unique headers, which may reflect browser-specific features or experimental headers. Firefox had a higher header similarity with Chrome than with Brave.
@@
Figure 2 represents a Venn diagram and provides a direct comparison of response headers between different browser versions, highlighting data and feature similarity. The intersection numbers, such as the 5672 headers common to all browsers, suggest a significant overlap in features across different platforms. However, the presence of headers unique to each browser or shared by only two browsers (e.g., 1427 unique to $\text{Chrome}_{22}$ and 154 unique to $\text{Brave}_{22}$) indicates that while there is a core set of features, there is also a notable amount of variation which can be critical for web tracker detection.
@@
The Empirical Cumulative Distribution Function plot for $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$ in Figure 3 row (A) show that the Content-Length values for non-trackers are generally higher than those for trackers, which suggests that non-trackers tend to send larger responses. This is consistent across both browser versions, indicating that while browsers evolve, certain characteristics like the smaller size of tracker responses remain consistent. The second row (B) plots the feature similarity for the X-XSS-Protection header, showing that header values can be very similar between the datasets, i.e., the training and test sets. It further supports our observation that the datasets are similar and that trackers and non-trackers have differences across the usage of HTTP header fields and values. Although we present only few selected examples here, they should highlight the general challenges that come with response messages datasets.
@@
Figure 3. Row (A) presents the ECDFs for both Chrome datasets for the Content-Length header values. We set a cut-off value at 10 000 to highlight the core observation and more than half of the responses have a value are below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$. Row (B) shows how similar the values across Chrome, Firefox, and Brave are for the X-XSS-Protection header.
Additional characteristics of the datasets include the high amount of missing values per header. One explanation might be that servers do not need to use every header field that exists, and it is common practice to use developer headers that are not part of the standard, i.e. they are not registered in the Request for Comments 4229111RFC 4229, https://www.ietf.org/rfc/rfc4229.txt, accessed: November 30, 2023. There are multiple headers that contain mostly unique values, e.g., Date. Lastly, according to the used filter lists, most HTTP response messages are not related to any tracking activities. This entails a class imbalance to be handled carefully to derive meaningful results from the final models.
@@ 5. Methodology
Based on our findings from the data exploration in Section 4, we propose our methodology to assess the effectiveness of classifiers based on response headers.
@@ 5.1. Problem Formulation
We consider the problem of binary classification, where we aim to predict the presence of web trackers based on response headers. The headers, represented by the feature vector $x\in\mathbb{R}^{d}$, are nominal and can be both discrete and continuous variables. These variables are derived from HTTP headers of four different browser types, denoted as the set $B=\{\text{Chrome},\text{Firefox},\text{Brave}\}$. For this study, we specifically consider versions $\text{Chrome}_{22},\text{Chrome}_{23},\allowbreak\text{Firefox}_{22},\text{Brave}_{22}$.
Given a training set $\mathcal{D}_{\text{train}}=\{(x_{1}y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ is a feature vector of the $i$-th HTTP header and $y_{i}\in\{0,1\}$ represents the associated class label (with ’1’ indicating the presence of a web tracker), our goal is to train a set of classifiers $f_{j}:\mathbb{R}^{d}\rightarrow\{0,1\}$ for $j\in M$ that minimizes the empirical risk:
@@
where $L(y_{i},f_{i}(x_{i}))$ is the loss incurred by predicting $f_{i}(x_{i})$ when the true label is $y_{i}$. The specific form of the loss function $L(\cdot)$ will be dependent on the particular classifier $f_{j}$ used.
We consider datasets represented by $D_{s}$, where $s$ can pertain to training, testing, validation, or calibration subsets of the data. We observe several challenges from our data exploration:
  1. (1)
    High Cardinality: The feature vectors $x_{i}$ in our datasets $D_{s}$ can exhibit high cardinality for certain features. This problem can pose challenges in terms of model training and encoding.
  2. (2)
    Sparseness: The feature vectors $x_{i}$ in our datasets $D_{s}$ are represented as sparse vectors, with each vector defined as:
@@
    Here, a significant proportion of the entries $v_{j}$ within $x_{i}$ are either zero or represent missing data, leading to the overall sparseness of $D_{s}$. Such sparseness can introduce computational complexities and may challenge the performance of certain machine learning algorithms not optimized for sparse data.
  3. (3)
    Class Imbalance: For datasets collected from a given browser $B$ and represented as $D_{s}$, the class distribution can exhibit imbalance. The majority class is denoted by $\alpha_{B}$, and the minority class is represented by $\beta_{B}$, satisfying $\alpha_{B}+\beta_{B}=1$. The degree of this imbalance can vary based on the browser as shown in Table 1, with some showing a more pronounced disparity between $\alpha_{B}\text{ and }\beta_{B}$. Such imbalance can introduce bias to models, favoring the majority class and potentially undermining the accurate prediction of the minority class.
  4. (4)
    High Dimensionality: With a high dimensionality $d$ and a high number of observations $n$ per dataset, the classification is susceptible to the Curse of Dimensionality, potentially leading to overfitting and necessitating intricate dimensionality reduction techniques. We have to apply a dimensionality reduction process that achieves a reduced feature space $X_{i}^{\prime}\subseteq X$, entailing a reasonable training time.
  5. (5)
    Data Similarity: The datasets $\mathcal{D}_{s}$ collected across different browsers can exhibit pronounced similarities in their feature distributions and values. Given the high dimensionality of our feature space, exhaustive pairwise comparisons can be computationally expensive and time-consuming, highlighting the challenge of accurately gauging data similarities in such a complex space.

### v2

@@ 3. Approach
This study addresses the research gap identified in our review of related research by evaluating the (cross-browser) performance of commonly used machine learning classifiers that are trained on HTTP response headers. In this section, we present our underlying research questions and the approach we employ to address them.
@@ 3.1. Preliminaries
A core principle of automated web tracker detection is delineating what constitutes a tracker. The focus of this delineation lies on the tracking entities and not on the tracking activity. Our definition of a tracker (T) and non-tracker (NT) is contingent upon the ground truth used to label each dataset:
@@
In other words: if the remote hostname (i.e., the sender) of a response matches at least one rule in our aggregate list compiled from the EL and EP filter lists, we label the response as a tracker.
For the sake of simplicity, we use the terms header and response header interchangeably to refer to HTTP response headers throughout the remainder of the paper unless explicitly stated otherwise.
@@ 3.2. Research Questions
The primary objective of this study is to assess the efficacy of response headers for automated web tracker detection. Based on the ML approaches of prior research and the observed areas to further extend our empirical understanding and detection of web trackers, we formulate the following research questions (RQ) that guide us throughout this study:
  • RQ1: What are the key characteristics of HTTP response header data in the web tracking domain? We want to understand the datasets to develop our pipeline and classifiers accordingly.
  • RQ2: Can trained classifiers detect web trackers with only response header features in an imbalanced setting while achieving a similar performance to past research? We aim to train high-performing classifiers while keeping the likely true distribution. We define our lower bound as an F1-Score of 0.9 based on the classification results of past research to provide a meaningful addition to existing detection solutions. Trackers and non-trackers are not equally present in the datasets, thus an imbalance is present and we want to assess how well classifiers perform in this setting.
  • RQ3: How large is a classifier’s performance degradation when it is trained on one browser’s HTTP traffic data and then applied to other browsers? Web trackers might exhibit different header characteristics depending on the browser. We use data from three browsers to capture a potential variety of web trackers and to assess the stability and generalization stability of our classifiers.
We address our RQs by developing a semi-automated ML pipeline that transforms HTTP traffic datasets collected with T.EX - The Transparency EXtension (T.EX) (Raschke et al., 2019) and use it to train several supervised binary classifiers. Additionally, we use reliable data compiled by prior research to facilitate future comparisons between new classifiers. Semi-automated means that the ML pipeline has to be started manually, is based on manual data exploration, and does not employ any AutoML libraries. The pipeline executes all steps automatically but cannot react to structurally new datasets without any changes or human intervention. Fully automating the pipeline requires the development of a direct interface between T.EX and the pipeline, i.e., automatically running new crawls and importing the new datasets to the pipeline, as well as structural converters for datasets from other sources.
To test the robustness and reliability of our classifiers, we train them on a Chrome dataset and apply them to other browser datasets (latitudinal comparison) and another Chrome dataset generated one year later (longitudinal comparison). Furthermore, as past research observed an uneven distribution of trackers and non-trackers, we train our classifiers with the original imbalanced distribution to assess how effectively our classifiers perform on the true distribution.
@@
Table 1. Overview of the collected datasets from (Raschke, 2023a), split by datasets as well as tracker (T) and non-trackers (NT).
@@ 3.3. Data Collection
We use existing datasets containing labeled HTTP messages by EL and EP. Following best practices suggested by (Demir et al., 2022), the crawls that generate the datasets should fulfill three criteria: (i) repeated crawls across multiple browsers, (ii) documentation of the geographical region as well as the technical environment, and (iii) selection of web sites is based on the Tranco ranking, which was proven to be less biased than the Alexa ranking (Pochat et al., 2018). An open-source dataset by Raschke that used T.EX fulfills all three criteria and is described as follows:
@@
After completing the crawls, each instance’s extension storage was extracted […]. Note: 2 of the 6 Firefox instances crashed during the crawl. Therefore, only 4 Firefox datasets are available (Raschke, 2023a).
To assess potential concept drift and resulting model degradation over time, we crawled one more dataset for Chrome in 2023 and made it open-source (Raschke, 2023b). We follow a similar methodology as described in (Raschke and Cory, 2022), using T.EX and visiting the landing pages with Chrome to explore the top 10K websites from the Tranco ranking as of March 29th, 2023.
Our reasons for only generating an additional dataset for Chrome are as follows: firstly, Chrome is the most widely used browser with a market share of 66.13% on desktops, whereas Firefox has 7.1%  [^3] https://gs.statcounter.com/browser-market-share/desktop/worldwide and Brave has less than 1%  [^4] https://kinsta.com/browser-market-share/. Based on this distribution, the Chrome dataset serves as the foundation for training to reflect and reach the majority, i.e., using data most users encounter. Secondly, web privacy research often relies on a simulated Firefox browser with Selenium due to the use of OpenWPM, leaving space for other browsers of interest. Thirdly, it helps us answer whether a model based on Chrome data can be applied to Firefox or Brave while achieving high-performance metrics.
@@ 4. Initial Response Header Analysis
To address RQ1 and gain a better understanding of the data underlying our analysis before designing our ML pipeline, we conduct an initial analysis of the data, the results of which we present in this section. Our exploration is guided by intuition and domain knowledge. We first explore headers across all three browsers to find empirical evidence of their potential relation to trackers. The goal is to identify the key characteristics of the datasets and not to present a complete analysis. We then use these insights to develop our classifiers and evaluate the effectiveness of response headers for web tracker detection.
@@ 4.1. Descriptive Analysis of Headers
First, we conduct a descriptive analysis to understand the data structure and distribution of class and feature features within the datasets. Table 1 summarizes descriptive metrics for each dataset. We begin with the datasets $\text{Chrome}_{22}$, $\text{Chrome}_{23}$, and $\text{Firefox}_{22}$. All three datasets show similar results for the distribution between trackers and non-trackers, as well as the number of unique headers in total and per response.
Let $B$ represent the set of browsers, where $B=\{\text{Chrome}_{22},\allowbreak\text{Chrome}_{23},\allowbreak\text{Firefox}_{22},\text{Brave}_{22}\}$ and let $L$ represent the set of labels, where $L=\{T,NT\}$. For each browser $b\in B$ and each label $l\in L$, we define $HN_{b,l}$ as the set of all distinct header names across all HTTP response messages from a browser $b$ and categorized under label $l$. The number of unique headers for each combination of $b$ and $l$ is defined as the cardinality of $HN_{b,l}$:
$|HN_{b,l}|=|\{hn\mid hn\in HN_{b,l}\}|$
We can observe a small quantitative change in the number of unique headers between $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$. On average, non-trackers have more unique headers as well as a higher median $\tilde{\mu}$ and interquartile range than trackers.
@@
As Brave is a privacy-oriented browser that performs extensive blocking of web tracking by default, $\text{Brave}_{22}$ exhibits a far lower number of trackers and unique headers in tracking-related response messages than the other datasets. Interestingly, the $\tilde{\mu}$ equals 18 for trackers and 15 for non-trackers, representing the only instance where trackers had a higher number. Although the dataset comprises almost exclusively non-trackers, the number of unique headers is similar to the other datasets.
Figure 1. Frequency distribution of headers for each browser, indicating the frequency with which each header is present in the dataset.
@@
Figure 1 shows a frequency plot to outline how often each header appears in the dataset. The plot demonstrates a power-law distribution, where few headers occur frequently, and many others form a long tail of low frequency. This sparseness indicates that while there are many different headers, only a few are consistently used across different websites and by different browsers.
@@
Figure 2. Venn diagram of HTTP response headers in the examined datasets, revealing a core set of 5672 headers that are common across Chrome, Firefox, and Brave. Unique headers are present in each browser, although $\text{Chrome}_{23}$ has the highest number of unique headers, which may reflect browser-specific features or experimental headers.
@@
Figure 2 depicts a Venn diagram that compares the response headers present in the examined datasets, highlighting data and feature similarity between the different browsers. The intersection numbers, such as the 5672 headers common to all browsers, suggest a significant overlap in features across different platforms. However, the presence of headers unique to each browser or shared by only two browsers indicates that while there is a core set of shared features, there is also a notable amount of variation that may be critical for web tracker detection.
@@
For instance, the Empirical Cumulative Distribution Function plots for $\text{Chrome}_{22}$ and $\text{Chrome}_{23}$ presented in Figure 3 row (A) show that the Content-Length values of non-trackers are generally higher than those of trackers, which suggests that non-trackers tend to send larger responses. This is consistent across both datasets, indicating that while browsers evolve, certain characteristics, like the smaller size of tracker responses, remain consistent. Conversely, row (B) of Figure 3 plots the feature similarity for the X-XSS-Protection header, showing that header values can be very similar between the datasets, i.e., the training and test sets. This further supports our observation that most of the examined datasets are similar and that trackers and non-trackers differ in their use of HTTP header fields and values. Although we have presented only a few selected examples here, they highlight the general challenges that come with response message datasets.
@@
Figure 3. Row (A) presents the ECDFs for both Chrome datasets for the Content-Length header values. We set a cut-off value at 10,000 to highlight the core observation and more than half of the responses have a value below this threshold – $\tilde{\mu}_{T22}=62$, $\tilde{\mu}_{NT22}=8068$ and $\tilde{\mu}_{T23}=45$, $\tilde{\mu}_{NT23}=8021.5$. Row (B) shows how similar the values across $\text{Chrome}_{22}$, $\text{Firefox}_{22}$, and $\text{Brave}_{22}$ are for the X-XSS-Protection header.
Additional characteristics of the datasets include a high amount of missing header values, resulting from a large number of available HTTP headers, of which any given response only uses a small number. This is exacerbated by the common practice of using custom headers that are not part of the HTTP standard, i.e., are not registered in RFC 4229555https://www.ietf.org/rfc/rfc4229.txt. This results in the presence of 8058 custom headers in the $\text{Chrome}_{22}$ dataset, contrasted by a comparatively small $\tilde{\mu}$ of unique headers per response as shown in Table 1.
The examined datasets are formatted as a matrix, with rows representing individual HTTP requests or responses and each unique HTTP header denoted in a separate column. This, combined with the observed large total number of unique headers and small number of headers per response, results in a sparse matrix where most elements are empty.
@@ 5. Methodology
Based on the findings of our data exploration presented in Section 4, we propose a comprehensive methodology to assess the effectiveness of classifiers based on response headers.
@@ 5.1. Problem Formulation
We consider the problem of binary classification, where we aim to predict the presence of web trackers based on response headers. The headers, represented by the feature vector $x\in\mathbb{R}^{d}$, are nominal and can be discrete and continuous variables. These variables are derived from HTTP headers of four different browser types, denoted as the previously defined set $B$. For this study, we specifically consider the datasets Chrome22, Chrome23, Firefox22, and Brave22.
Given a training set $\mathcal{D}_{\text{train}}=\{(x_{1}y_{1}),\ldots,(x_{n},y_{n})\}$, where $x_{i}$ is a feature vector of the $i$-th HTTP header and $y_{i}\in\{0,1\}$ represents the associated class label, our goal is to train a set of classifiers $f_{j}:\mathbb{R}^{d}\rightarrow\{0,1\}$ for $j\in M$ that minimizes the empirical risk:
@@
where $L(y_{i},f_{i}(x_{i}))$ is the loss incurred by predicting $f_{i}(x_{i})$ when the true label is $y_{i}$. The specific form of the loss function $L(\cdot)$ will be dependent on the classifier $f_{j}$ used.
We consider datasets represented by $D_{s}$, where $s$ can pertain to training, testing, or calibration subsets of the data. We observe several challenges (CH) and key characteristics from our data exploration, which answer our RQ1:
  • (CH1) High Cardinality: The feature vectors $x_{i}$ in our datasets $D_{s}$ may exhibit high cardinality for certain features. This problem may pose challenges in terms of model training and encoding.
  • (CH2) Sparseness: As discussed above, the feature vectors $x_{i}$ in our datasets $D_{s}$ are represented as sparse vectors, with each vector defined as:
@@
    A significant proportion of the entries $v_{j}$ within $x_{i}$ are either zero or represent missing data, resulting in an overall sparseness of $D_{s}$. Such sparseness may introduce computational complexities and challenge the performance of certain ML algorithms that are not optimized for sparse data.
  • (CH3) Class Imbalance: For datasets collected from a given browser $B$ and represented as $D_{s}$, the class distribution may be imbalanced. The majority class is denoted by $\alpha_{B}$, and the minority class is represented by $\beta_{B}$, satisfying $\alpha_{B}+\beta_{B}=1$. The degree of this imbalance may vary based on the browser, as shown in Table 1, with some datasets exhibiting a more pronounced disparity between $\alpha_{B}\text{ and }\beta_{B}$. Such imbalance can introduce bias to models, favoring the majority class and potentially undermining the accurate prediction of the minority class.
  • (CH4) High Dimensionality: With a high dimensionality $d$ and a high number of observations $n$ per dataset, the classification is susceptible to the Curse of Dimensionality, potentially leading to overfitting and necessitating intricate dimensionality reduction techniques. Therefore, applying a process that achieves a reduced feature space $X_{i}^{\prime}\subseteq X$ is necessary, entailing a reasonable training time.
  • (CH5) Data Similarity: The datasets collected across browsers can exhibit pronounced similarities in their feature distributions and values. Given the high dimensionality of our feature space, exhaustive pairwise comparisons can be computationally expensive and time-consuming, highlighting the challenge of accurately gauging data similarities in such a complex space.